OpenStudy (anonymous):

How do i find where the particle is at rest? The derivative is t(-1/5e^(-t/5))+e^(-t/5). I know to set it equal to zero but i cant figure out how to get the point

OpenStudy (anonymous):

k so u r looking for solutions for t so ull have two parts one is where t = 0 and the second will be at e^(-t/5)=e^(-t/5)/5

OpenStudy (anonymous):

then take ln of both sides and u get ln(e^(-t/5)) = ln(5(e^(-t/5)))

OpenStudy (anonymous):

lol m solving it with u so going step by step hope it helps

OpenStudy (anonymous):

there is only one set of points. one (x,y). how do i set this up? ahhh okay i'm seeing it i think... keep it up i'm understanding you lol

OpenStudy (anonymous):

is it ln of 5 + ln of e^(-t/5)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so i have -t/5= ln 5 -t/5....

OpenStudy (anonymous):

yea no tht doesnt work double check the integral ... was there a substitution u used?

OpenStudy (anonymous):

sorry i mean the derivative

OpenStudy (anonymous):

no i actually already put the derivative into the webassign and it was right... now it wants when the particle is at rest